# Electrical networks and data analysis in phylogenetics

### v1

@@ 3.2. Circular split systems and electrical networks
The pseudo metric associated to a spit $S$ is defined by the following matrix $D:$
$D_{S}(k,l)=\begin{cases}1,\ \text{if}\ |A\cap\{k,l\}|=1,\\ 0,\ \text{otherwise.}\end{cases}$
@@
We call a weighted circular split system a circular split system with a positive weight attached to each set of parallel edges which display the splits. Such a weighted circular split system defines a metric on the set $X$ by the formula
@@
The split ($(\{1,2,3,4\}|\{5,6,7,8\})$ in the Figure 4 corresponds to the set of edges with the weight $\alpha_{1}$.
Figure 4. Circular split system and their polygon representations
@@
Since the resistance metric defined by a planar circular electrical network satisfies the Kalmanson condition it corresponds to a unique weighted circular split system, which we call an electrical circular split system following [15] or, given the theorem 3.8, an electrical Kalmanson metric.
@@
The following theorem gives the formula for the weights of a weighted circular split systems 3.12.
@@
$\omega_{ij}=\frac{1}{2}(d_{i,j}+d_{i+1,j+1}-d_{i,j+1}-d_{i+1,j})$
@@
Here are our main results, they give a complete characterization of the Kalmanson metrics which are the resistance metrics of a circular electrical networks.
@@
Let $D$ be a Kalmanson metric matrix, then $D$ is the effective resistance matrix of a connected circular electrical network $\mathcal{E}$ if and only if the matrix $\Omega_{D}$ constructed from $D$ according to the formula (3.3) defines a point $X$ in $\mathrm{Gr}_{\geq 0}(n-1,2n)$ and the Pluecker coordinate $\Delta_{24\dots 2n-2}(X)$ does not vanish.
@@
It remains to prove that $d_{ij}$ are equal to the effective resistances $R_{ij}$ of the network $\mathcal{E},$ associated with $\Omega_{D}.$ Since the metrics $d_{ij}$ and $R_{ij}$ are both Kalmanson and the weights in their split decomposition are equal, $\omega_{ij}=-m_{ij}$ (see Formula (3.9)) , we conclude that:
@@
If any $k\times k$ circular minors of the matrix $M(D)$ is positive after multiplying by $(-1)^{k}$ then $M(D)$ defines a response matrix of a network $\mathcal{E}^{\prime}$ due to Theorem 2.4. Hence we conclude that $\mathcal{E}^{\prime}$ is connected 2.4 given the condition on the rank of $M(D)$.
@@
The resistance metric has the following important property its square root $\sqrt{D(i,j)}$ is $L_{2}$ embeddable [22], hence the electric Kalmanson metrics have that property. There is a well known condition for $L_{2}$ embeddability stated in terms of the minors of the Caley-Menger matrix [11]. It is interesting to see if this condition along describes the electric Kalmanson metrics in the set of all Kalmanson metrics.
@@
The formula (3.9) for $R_{ij}$ can be obtained from Theorem 3.6 since $\mathcal{E}^{**}=\mathcal{E}^{\prime}$, where $\mathcal{E}^{\prime}$ is obtained from $\mathcal{E}$ by shifting the numeration of the nodes of $\mathcal{E}$ by $1$ clockwise:
where $\{1,\dots,n\}$ and $\{\overline{1},\dots,\overline{n}\}$ are the labels of the nodes of $\mathcal{E}$ and $\mathcal{E}^{*}$ respectively.
Figure 5. Labeling in the formula (3.11)
@@
Many statements in this paper can be extended to the connected cactus networks. In particular, one can obtain that the effective resistances of a connected cactus network give rise to a Kalmanson pseudometric and a Kalmanson pseudometric matrix $D$ can be identified with the resistance matrix of a connected cactus networks if and only if the conditions similar to the conditions in Theorem 3.16 hold.
It would allow to conclude that for a Kalmanson metric the non vanishing of a certain Pluecker coordinate condition in Theorem 3.16 and the rank condition in Theorem 3.17 are automatically satisfied.
@@
Let $X$ and $D$ be as in Theorem 3.14, then the matrix $M(D)$ defined above is the response matrix of a connected electrical network, not necessarily circular. Namely it is symmetric, all non diagonal elements are negative and the sum of the elements in any row is zero. The connection of the resistance matrix of this network to the original metric $D$ is an interesting question.
Moreover such a matrix defines a point
$\mathrm{IG}(n-1,2n)\subset\mathrm{Gr}(n-1,2n)$
according to 3.7. Its Pluecker coordinates are interesting invariants of the metric $D$. We are planning to address these questions in a future publication.
@@
Let $T$ be a tree with four leafs as in the Figure 6 with the weights of edges all equal to one. The resistance metric matrix of the tree is
@@ 4. Reconstruction of network topology
[8], [17] If a matrix $D$ is tree realizable, then there is an unique minimal tree $T_{min}$ such that $D=D_{T_{min}}$, where the minimality of $T_{min}$ means that for any other $T$ with the property $D=D_{T}$, $|E(T_{min})|<|E(T)|.$
@@
The median graph of a circular network $\mathcal{E}$ with the graph $\Gamma$ is the graph $\Gamma_{M}$ whose internal vertices are the midpoints of the edges of $\Gamma$ and two internal vertices are connected by an edge if the edges of the original graph $\Gamma$ are adjacent. The boundary vertices of $\Gamma_{M}$ are defined as the intersection of the natural extensions of the edges of $\Gamma_{M}$ with the boundary circle. Since the interior vertices of the median graph have degree four, we can define the strands of the median graph as the paths which always go straight through any degree four vertex.
@@
Denote by $A_{i}$ the columns of the matrix $\Omega_{R}(\mathcal{E})$ and define the column permutation $g(\mathcal{E})$ as follows: $g(\mathcal{E})(i)=j,$ if $j$ is the minimal number such that $A_{i}\in\ \mathrm{span}(A_{i+1},\dots,A_{j}),$ where the indexes are taking modulo $2n$.
@@
This result follows from the fact that $\Omega_{R}(\mathcal{E})$ is an explicit parametrization of the Lam embedding, see Section $4.6$ [24]. ∎
@@
Based on Theorem 4.6 we suggest the following reconstruction algorithm:
@@
If there is a tree $T$ among the graphs representing a minimal circular electrical network $\mathcal{E}\in E_{n}$ then it is unique.
@@
Indeed, it is easy to see that $R_{T}$ is a tree realizable by both $\bar{T}$ and $T_{min},$ therefore $|E(T)|\geq|E(T_{min})|.$ Since $T$ and $\bar{T}_{min}$ have the same effective resistant matrix, they are equivalent. Therefore due to Theorem 4.8 $\bar{T}_{min}$ can be transform to $T$ which implies $|E(T_{min})|\geq|E(T)|$ which contradicts to the uniqueness of the minimal tree. ∎
@@
  • Transform $\mathcal{E}$ to a minimal tree by a sequence of the star-triangle transformation. By Theorem $1$ from [3] we suppose that heuristically this converting can be done monotonically by changing all the triangles to the stars.
@@
A more advanced technique called the chamber ansatz when applied to $\Omega_{R}(\mathcal{E})$ gives an algorithm for recovering not only the topology of the network but the weights of the edges as well. We will describe this method in a coming work [19].
@@
Then the matrix $\Omega^{\prime}_{R}(\mathcal{E})\in Gr_{\geq 0}(n-1,2n)$ has the form
$\Omega^{\prime}_{R}(\mathcal{E})=\begin{pmatrix}1&3&1&1&0&-1&0&1\\ 0&1&1&2&1&1&0&0\\ 0&-1&0&1&1&3&1&1\end{pmatrix}$

### v2

@@ 3.2. Circular split systems and electrical networks
The pseudo metric associated with a split $S$ is defined by the following matrix $D:$
$D_{A|B}(k,l)=\begin{cases}1,\ \text{if}\ |A\cap\{k,l\}|=1,\\ 0,\ \text{otherwise.}\end{cases}$
@@
We call a weighted circular split system a circular split system with a positive weight attached to each set of parallel edges, which display the splits. Such a weighted circular split system defines a pseudo metric on the set $X$ by the formula
@@
The split ($(\{1,2,3,4\}|\{5,6,7,8\})$ of the set of the sides of the polygon on the left defined by the yellow diagonal in the Figure 4 corresponds to the set of parallel edges with the weight $\alpha_{1}$ in the graph on the left.
Figure 4. Circular split system and their polygon representations. Removing any set of parallel edges defines a split of the set of leaves of the graph on the right which coincides with the split of the set of the sides of a polygon on the left obtained by removing an appropriate diagonal.
@@
Since the resistance metric defined by a planar circular electrical network satisfies the Kalmanson condition, it corresponds to a unique weighted circular split system, which we call an electrical circular split system following [15] or, given Theorem 3.9, an electrical Kalmanson metric.
@@
The following theorem gives the formula for the weights of a weighted circular split systems 3.14.
@@
$\omega_{ij}=\frac{1}{2}(d_{i,j}+d_{i+1,j+1}-d_{i,j+1}-d_{i+1,j}).$
Remark 3.17.
The formula (3.9) for $R_{ij}$ can be obtained from Theorem 3.7 since taking the dual electrical network is an involution up to a shift: $\mathcal{E}^{**}=\mathcal{E}^{\prime}$, where $\mathcal{E}^{\prime}$ is obtained from $\mathcal{E}$ by shifting the numeration of the nodes of $\mathcal{E}$ by $1$ clockwise:
where $\{1,\dots,n\}$ and $\{\overline{1},\dots,\overline{n}\}$ are the labels of the nodes of $\mathcal{E}$ and $\mathcal{E}^{*}$ respectively. Thus Theorem 3.7 in fact provides an alternative way to observe that the effective resistance distance is the Kalmanson metric.
Figure 5. Labeling in the formula (3.10)
@@
Given a Kalmanson metric as a matrix $D$, one can decide whether it is an electrical Kalmanson metric in a straightforward but computationally heavy way. First one must reorder the matrix columns and rows to ensure its entries satisfy the Kalmanson inequalities, then find the unique response matrix $M$ that corresponds to a given Kalmanson metric matrix using the pseudoinverse formula, see [15] (Lemma 3.1), and then check all the circular minors of that $M$ for non-negativity.
Here is our main result, it gives a new characterization of the Kalmanson metrics, which are the resistance metrics of circular electrical networks without using inversion of matrices.
@@
Let $D$ be a Kalmanson metric matrix with the correct order of rows and columns, then $D$ is the effective resistance matrix of a connected circular electrical network $\mathcal{E}$ if and only if the matrix $\Omega_{D}$ constructed from $D$ according to the formula (3.3) defines a point $X$ in $\mathrm{Gr}_{\geq 0}(n-1,2n)$ and the Plücker coordinate $\Delta_{24\dots 2n-2}(X)$ does not vanish.
@@
It remains to prove that $d_{ij}$ are equal to the effective resistances $R_{ij}$ of the network $\mathcal{E},$ associated with $\Omega_{D}.$ Since the metrics $d_{ij}$ and $R_{ij}$ are both Kalmanson and the weights in their split decomposition are equal, $\omega_{ij}=-m_{ij}$ (see the formula (3.9)) , we conclude that:
@@
The following theorem is also a convenient test for detecting electrical Kalmanson metrics. It literally falls in our lap due to Theorems 3.7 and 3.16.
@@
If any $k\times k$ circular minors of the matrix $M(D)$ is positive after multiplying by $(-1)^{k}$ then $M(D)$ defines a response matrix of a network $\mathcal{E}^{\prime}$ due to Theorem 2.4. Hence we conclude that $\mathcal{E}^{\prime}$ is connected (see Theorem 2.4) given the condition on the rank of $M(D)$.
@@
Many statements in this paper can be extended to the connected cactus networks. In particular, one can obtain that the effective resistances of a connected cactus network give rise to a Kalmanson pseudometric and a Kalmanson pseudometric matrix $D$ can be identified with the resistance matrix of a connected cactus networks if and only if the matrix $\Omega_{D}$ constructed from $D$ according to the formula (3.3) defines a point $X$ in $\mathrm{Gr}_{\geq 0}(n-1,2n)$.
@@
In Remark 4.3, [F] S. Forcey gives an example of a non-planar network whose resistance metric is nevertheless Kalmanson. It is easy to generalise this example in the following way. Let’s start with a circular planar electrical network, whose resistance metric therefore is Kalmanson, such that all the inequalities required by the Kalmanson condition are strict. This means the resistance metric is represented by a full circular planar split system. Add an edge with a variable conductance $x$ to it in such a way that it is not a circular planar anymore, as in the Figure 6. It is clear from the Kirchhoff formula 3.2 that each effective resistance $R_{ij}$ changes as a function of $x$ according to the formula
$R^{\text{n}ew}_{ij}=\frac{R_{ij}+r_{ij}x}{1+Tx}$
where $r_{ij}$ and $T$ are non-negative numbers and $T$ does not depend on $i,\,j$. Therefore, for small $x$ the Kalmanson condition for the resistance metric will hold for the new network, although it is not any longer circular planar. This means that the planar circular Kalmanson metrics are sitting far away from the boundary of the space of all Kalmanson metrics.
Figure 6. A non-planar electrical network with nodes {1,…,5}
@@
Let $X$ and $D$ be as in Theorem 3.16, then the matrix $M(D)$ defined above is the response matrix of a connected electrical network, not necessarily circular planar. Namely, it is symmetric, all non-diagonal elements are negative and the sum of the elements in any row is zero. The connection of the resistance matrix of this network to the original metric $D$ is an interesting question.
Moreover, such a matrix defines a point
$\mathrm{IG}(n-1,2n)\subset\mathrm{Gr}(n-1,2n)$
according to Theorem 3.8. Its Plücker coordinates are interesting invariants of the metric $D$. We are planning to address these questions in a future publication.
@@
The resistance metric has the following important property: its square root $\sqrt{D(i,j)}$ is $L_{2}$ embeddable [23], hence the electric Kalmanson metrics have that property. There is a well known condition for $L_{2}$ embeddability stated in terms of the minors of the Caley-Menger matrix [11]. It is interesting to see if this condition also describes the electric Kalmanson metrics in the set of all Kalmanson metrics.
@@
Let $T$ be a tree with four leaves as in the Figure 7 with the weights of edges all equal to one. The resistance metric matrix of the tree is
@@ 4. Reconstruction of network topology
[8], [17] If a matrix $D$ is tree realizable, then there is an unique tree $T$ such that $D=D_{T}.$
@@
The median graph of a circular network $\mathcal{E}$ with the graph $\Gamma$ is the graph $\Gamma_{M}$ whose internal vertices are the midpoints of the edges of $\Gamma$ and two internal vertices are connected by an edge if the edges of the original graph $\Gamma$ are adjacent in one face. The boundary vertices of $\Gamma_{M}$ are defined as the intersection of the natural extensions of the edges of $\Gamma_{M}$ with the boundary circle. Since the interior vertices of the median graph have degree four, we can define the strands of the median graph as the paths which always go straight through any degree four vertex.
@@
Denote by $A_{i}$ the columns of the matrix $\Omega_{R}(\mathcal{E})$ and define the column permutation $g(\mathcal{E})$ as follows: $g(\mathcal{E})(i)=j,$ if $j$ is the minimal number such that $A_{i}\in\ \mathrm{span}(A_{i+1},\dots,A_{j}),$ where the indices are taken modulo $2n$.
@@
This result follows from the fact that $\Omega_{R}(\mathcal{E})$ is an explicit parametrization of the Lam embedding, see Proposition $5.17$ [25]. ∎
@@
Based on Theorem 4.6 we suggest the following reconstruction algorithm (see Example 4.11):
@@
Let $\mathcal{E}$ be a minimal electrical network whose graph is a tree $T$, then such a tree is unique.
Let $\mathcal{E}(T,\omega)$ be a circular planar electrical network on a tree $T$ such that all its leaves are nodes and there are no inner vertices of degree $2$, then $\mathcal{E}(T,\omega)$ is minimal.
@@
Indeed, it is easy to see that $R_{T}$ is a tree realizable by $\bar{T}$. Therefore, $T$ is unique by Theorem 4.3. This proves the first statement.
To prove the second statement, it is enough to verify that the median graph $T_{M}$ of $\mathcal{E}(T,\omega)$ has no lenses. Indeed, each strand of $T_{M}$ corresponds to a simple path between nodes in the original graph, see the Figure 10. Since $T$ has no inner vertices of degree $2$ these paths can intersect at most once. Therefore, $T_{M}$ has no lenses and $\mathcal{E}(T,\omega)$ is minimal. ∎
Figure 10. On the left: A path which corresponds to a strand; On the right: the stands around a node of the degree $2$
@@
  • Transform $\mathcal{E}$ to a minimal tree by a sequence of the star-triangle transformation.
Heuristically, by ([3], Theorem $1$), we assume that the second step can be performed monotonically, converting all triangles into stars.
@@
A more advanced technique called the chamber ansatz when applied to $\Omega_{R}(\mathcal{E})$ gives an algorithm for recovering not only the topology of the network but the weights of the edges as well. This method is described in [19].
@@
Then the matrix $\Omega^{\prime}_{D}$ has the form
$\Omega^{\prime}_{D}=\begin{pmatrix}1&3&1&1&0&-1&0&1\\ 0&1&1&2&1&1&0&0\\ 0&-1&0&1&1&3&1&1\end{pmatrix}$
and satisfies the Theorem 3.19, therefore $\Omega^{\prime}_{D}=\Omega^{\prime}_{R}(\mathcal{E})$ for a network $\mathcal{E}\in E_{4}.$